OpenStudy (anonymous):

The sum of two numbers is 124. Two times the greater exceeds three times the smaller by 84. Find the numbers.

OpenStudy (anonymous):

can anybody help with this problem

OpenStudy (niel):

124=2n+.... hold on

myininaya (myininaya):

x+y=124 assume x is the greater number 2x=3y+84

myininaya (myininaya):

you have 2 equations you should be able to find x and y

OpenStudy (niel):

well nvm its pretty simple from there... just plug in 0 for x and then do the math to find y and do the same for y to find x

myininaya (myininaya):

x+y=124 means y=124-x so we can put the second equation in terms of x 2x=3(124-x)+84 2x=372-3x+84 2x+3x=456 5x=456 x=456/5 so y=124-x=124-456/5=164/5

OpenStudy (niel):

myininaya did an awesome job at getting that out fast... i had to try and remember the concept

OpenStudy (anonymous):

thank you myininaya and Niel